# Confronting Cold New Early Dark Energy and its Equation of State with Updated CMB and Supernovae Data

### v1

@@ III.2 Confronting NEDE with NPIPE (and Pantheon+)
We find that Cold NEDE reduces the $Q_{\mathrm{DMAP}}$ tension from $6.3\sigma$ within $\Lambda$CDM to $2.8\sigma$ (Plik/Planck 2018), $3.5\sigma$ (CamSpec/NPIPE), and $3.7\sigma$ (HiLLiPoP/NPIPE). When comparing this with the previous works mentioned above, the residual tension with the newer datasets is larger by up to $1.5\sigma$. However, this comparison has to be taken with a grain of salt: The previous analyses, which relied on Pantheon and Plik/Planck 2018, also gave rise to a weaker $\Lambda$CDM tension. For example, [19] (Plik/Planck 2018+BAO+Pantheon) used a prior on $H_{0}$ that corresponded to a $\Lambda$CDM $Q_{\mathrm{DMAP}}$ tension of $4.7\sigma$, which was reduced to $1.4\sigma$ within Cold NEDE. In other words, the level of relative improvement is similar to what we find here for Plik/Planck 2018 (in both cases, the relaxation of the tension is larger than $3\sigma$). Exchanging Plik/Planck 2018 with CamSpec/NPIPE diminishes this improvement slightly to around $2.5\sigma$. In all cases, we find a significant $\chi^{2}$ improvement of at least 28 units. While compared to previous analyses, the base model remains similarly efficient at reducing the Hubble tension, it cannot fully resolve it anymore (provided we accept that the Hubble tension has passed the $Q_{\mathrm{DMAP}}$ threshold of $6\sigma$). The observation that NPIPE data slightly disfavors the NEDE model compared to Plik/Planck 2018 is also supported by the contours in Fig. 1. Comparing the blue (Plik/Planck 2018) with the green (CamSpec/NPIPE) contours, it is apparent that the evidence for NEDE is diminished. To be specific, we find $f_{\mathrm{NEDE}}<8.3\%$ ($95\%$ C.L.) for NPIPE and $f_{\mathrm{NEDE}}<11.9\%$ ($95\%$ C.L.) for Planck 2018.555We stress that this should be understood as a relative comparison as both values are known to be affected by prior volume effects that become sizeable in the $f_{\mathrm{NEDE}}\to 0$ limit [18, 62]. A profile likelihood for $f_{\mathrm{NEDE}}$ can be expected to support a non-vanishing fraction of $f_{\mathrm{NEDE}}$ [58, 24]. Moreover, it should be noted that we are comparing $f_{\mathrm{NEDE}}$ at the time of decay, which is not exactly the same in Plik and NPIPE fits. Specifically, the later decay found for NPIPE might in part explain the difference. This result is reproduced by the orange contours (HiLLiPoP/NPIPE). Overall, we find good agreement between the CamSpec and HiLLiPoP likelihood code. [^6] In addition, we also investigated the internal consistency of the CamSpec likelihood within NEDE by combining our baseline datasets with only the NPIPE TT and the TEEE data. We find consistency between these reduced NPIPE runs and the full one. In particular, we find that polarization data alone has limited constraining power, leaving, for example, $w_{\mathrm{NEDE}}$ and $f_{\mathrm{NEDE}}$ rather unconstrained. The only exception, albeit not significant, concerns the constraint on trigger dark matter $\Omega_{\phi}$, which is $\Omega_{\phi}<0.8\%$ ($95\%$ C.L.) for HiLLiPoP and strengthens to $\Omega_{\phi}<0.5\%$ for CamSpec. In that context, it is important to note that in both cases the model preserves its ability to reduce the value of $S_{8}$, thereby keeping the Gaussian $S_{8}$ tension below $2.1\sigma$ for all analyses performed here (using the weak lensing constraint in [56] as a reference value). Moreover, we find in agreement with the previous analysis in [19] that including the SH0ES prior further reduces the tension to below $2\sigma$. This trend is particularly pronounced in the case of HiLLiPoP.
@@
In Fig. 2, we compare Cold NEDE (in blue) and AxiEDE (in red). We find that both models are similarly affected by NPIPE data, although the constraints on $f_{\mathrm{NEDE}}$ (and hence $H_{0}$) are somewhat weaker, leading to a marginally smaller residual tension in the case of Cold NEDE (in [26] a residual $3.7\sigma$ tension was reported for CamSpec). This raises the expectation that newer CMB datasets will be increasingly capable of distinguishing between the two models (see also Refs. [63, 11] for other comparisons). As mentioned before, while both models share similarities on background level, differences arise predominantly on perturbation level. To be specific, the perturbations in AxiEDE correspond to perturbations of a homogeneous scalar field background and are described by the perturbed Klein-Gordon equation, whereas perturbations in Cold NEDE, are perturbations of a barotropic fluid with an amplitude that, on sub-horizon-scales, is controlled by the momentum-dependent perturbations of a subdominant scalar field evaluated at $\tau_{*}$.
@@ III.3 Constraining a Time-Dependent Equation of State
Figure 3: Left: Comparison of the 2D posterior distributions obtained for the extended NEDE model with a time-dependent equation of state using either the CamSpec/NPIPE (red) or the Plik/Planck 2018 (blue) likelihood (alongside our baseline datasets and the SH0ES prior on $M_{b}$). The green curve shows the prediction for the two-fluid mixture in (17) and (18) (stiff fluid and radiation). The gray lines in the $w$ vs $dw/d\ln a$ plane show the case when the stiff fluid decays into radiation, where the dashed, dash-dotted and dotted lines correspond to $Q_{*}=1/2$, $Q_{*}=1$ and $Q_{*}=2$, respectively. Right: The distribution for the first two derivatives of the equation of state parameter restricted to $w_{\mathrm{NEDE}}(a_{*})=5/6$, demonstrating partial insensitivity to the value of the second derivative (in the case of CamSpec/NPIPE). The black line corresponds to $y=0$. In the region below it, $w$ would turn negative at a later time.
@@
In the $w$ vs $dw/d\ln a$ plane of Fig. 3, we observe an approximate degeneracy between these two parameters. A larger initial value of $w$ can be compensated by a more negative derivative. A similar degeneracy is observed in the other two panels that depict the $w$ vs $d^{2}w/d(\ln a)^{2}$ and the $dw/d\ln a$ vs $d^{2}w/d(\ln a)^{2}$ planes. There, however, one finds that the width of the degeneracy band increases towards large values of $w$. This feature reflects the fact that in that parameter region, the second derivative is less important for the parametrization of $w(a)$ and does not impact cosmological observables much. At the same time, there is a relatively rigid bound from below corresponding to $d^{2}w/d(\ln a)^{2}\gtrsim 0$ for both data set combinations. This, however, is due to the fact that in our simplified parametrization (10), that parameter region would lead to negative $w(a)$ at late times which is excluded by construction. This is demonstrated in the right panel of Fig. 3 for the fiducial value $w_{\mathrm{NEDE}}(a_{*})=5/6$. In the region below the red line, $w$ turns negative before today. We reiterate that this is an artifact of the truncation in (10). In particular, including higher-order terms is expected to relax that bound. This, however, is not expected to change the results for $dw/d\ln a$, as data can only constrain $w(a)$ over a limited redshift range close to $z_{*}$.
@@
The green curves show the two-fluid constraints (17) and (18) for a mixture with $w_{1}=1$ and $w_{2}=1/3$. Each point on the curve corresponds to a certain choice of the initial ratio $r$ defined in (16). These results suggests that the mixture model is consistent with both dataset combinations when the two components initially have similar energy densities or the stiff component dominates. For example, for the case of the Plik/ Planck 2018 likelihood, the preferred values of $w_{*}$ translate into the fluid ratio $r_{f}=0.68^{+0.20}_{-0.13}$ ($68\%$ C.L.), according to (16). The gray curves in the figure show the case when the stiff fluid decays into the radiation component with a decay rate controlled by $Q$, according to (20). The dashed, dash-dotted and dotted lines correspond to $Q_{*}=1/2$, $Q_{*}=1$ and $Q_{*}=2$, respectively. For these order unity values, a sizeable fraction of the stiff fluid decays within one Hubble time. In particular, we see that the decay scenarios are compatible with data as long as the decay rate satisfies $Q*\lesssim 2$. In particular, data appears to be compatible with a scenario where $Q_{*}=2$ and $r=1$, i.e., all of the latent heat is initially converted into a stiff fluid before it decays into radiation rather quickly. The gray curves are not shown in the $d^{2}w/dlna^{2}$ planes because they also depend on the time-dependence of the decay rate, i.e. on the parameter $q_{*}$ according to (22) (making these contours less predictive for the decay scenario).
@@
There are two takeaways from this analysis. First, the time-evolving equation of state is only (mildly) favored for the Planck 2018 data, while it is less constrained for NPIPE data. For both NPIPE and Planck 2018 data, there is a clear (and expected) degeneracy between $dw/d(\ln a)$ and $w^{*}$. Moreover, both datasets are fully compatible with the assumption $dw/d(\ln a)=0$ made in the base model, and accordingly the data fit only gives a very weak evidence for the extension. Second, in both cases, data appears to be compatible with a two-fluid system with $w_{1}=1$ and $w_{2}=1/3$, including the possibility of an energy transfer from the stiff to the radiation fluid. This motivates a more detailed study where both fluid components including their perturbations are evolved independently, without relying on the truncation.
@@ III.4 A First Test of the Thawing Assumption
Remaining agnostic about the underlying model, here, we treat $H_{*}/m$ as a free parameter and fit it to both the CamSpec/NPIPE and Plik/Planck 2018 likelihoods (together with our baseline datasets and the SH0ES prior on $M_{b}$). To be precise, we consider the model with parameters $\log_{10}(z_{*})$, $f_{\mathrm{NEDE}}$, $w_{*}$ ($=\mathrm{const}$), and $H_{*}/m$ (while we set $\Omega_{\phi}\ll 1$ for simplicity). Note that on the background level, for a fixed $z_{*}$, this parameter controls the value of $m$, which enters the equations of motion (4) for the trigger field. Since the trigger field is usually very subdominant, changing its mass has negligible effect on the cosmic evolution. At the perturbation level, it however affects the matching equations (7) through their dependence on $\delta\phi$ and $\bar{\phi}^{\prime}$. In general, they diverge for $\bar{\phi}^{\prime}\to 0$, which during radiation domination imposes the lower bound $H_{*}/m>0.12$. On the other hand, for $H_{*}/m\gg 1$, we enter the slow-roll regime, where the trigger is assumed to be inactive. We impose the flat prior $10.18<H_{*}/m<$ 1. The lower bound corresponds to the first zero of $\bar{\phi}$ after it has dropped out of slow-roll. In a specific trigger scenario such as Cold NEDE, we thus expect the phase transition to occur earlier, i.e., for larger values of $H_{*}/m$. The upper bound corresponds to a situation where the transition is triggered while $\phi$ is still marginally within the slow-roll regime (for even higher values the trigger perturbations are too suppressed for data to be sensitive to differences between different values of $H_{*}/m$).
@@ IV Conclusion
The Cold NEDE model introduces an early dark energy component that decays in a phase transition triggered by an ultralight scalar field. In the past, the model has been found to reduce both the $H_{0}$ and $S_{8}$ tensions below the $2\sigma$ level. It admits a simple field theoretic realization in terms of a first-order vacuum phase transition that occurs in the dark sector before matter-radiation equality. One of the benefits of the Cold NEDE model is its predictivity as it directly affects the physics of waves in the primordial plasma at scales that can be observed in the CMB and matter power spectra. This happens at the background level through a change of the sound horizon and at the perturbation level through the gravitational impact of dark sector perturbations associated with the decaying early dark energy fluid. Nevertheless, the model comes with freedom in how to realize the dissipation of the early dark energy component after the phase transition, which so far has been modeled in terms of an ideal fluid with a constant equation of state parameter $w_{\mathrm{NEDE}}$. At the same time, new CMB, LSS and supernovae observations offer the prospect of further constraining early dark energy models and, in the case of Cold NEDE, this should enable us to further constrain $w_{\mathrm{NEDE}}$ after the phase transition, providing a way to further narrow down viable microscopic scenarios.
@@
In line with this overarching goal, in this work, we have analyzed the cold NEDE model and some of its physically well motivated extensions in light of updated CMB and supernovae data. To be precise, the $-\ell$ Planck 2018 polarization and temperature data and the Pantheon supernovae data were replaced by Planck NPIPE (using the CamSpec and HiLLiPoP likelihoods) and Pantheon+ data, respectively. While we find that the model’s ability to resolve the Hubble tension is somewhat diminished with the inclusion of the new datasets, it still reduces the $Q_{\mathrm{DMAP}}$ tension considerably from $6.3\sigma$ within $\Lambda$CDM to $3.5\sigma$ when using the CamSpec likelihood code for the NPIPE data (and $3.7\sigma$ for HiLLiPoP). This outcome is compatible with recent findings for the AxiEDE model, where a residual $3.7\sigma$ tension was reported [26]. It is important to note that the improvement over $\Lambda$CDM is almost unchanged compared to earlier works when it comes to relieving the Hubble tension, and the increased residual tension is mostly due to $\Lambda$CDM itself showing a stronger tension when using NPIPE and Pantheon+ data. It is also worth noting that the new DESI BAO data, that we have left out of this analysis, would likely reduce the tension level given results for the axiEDE model [51, 52]. Moreover, we find that the model’s ability to help with the $S_{8}$ tension is maintained in the presence of the new datasets. Most notably, in contrast to other EDE type models, including the SH0ES prior lowers the value of $S_{8}$ due to the presence of the trigger oscillations that act as a small fuzzy dark matter component at late time [19]. For concreteness, when using the weak lensing constraint from [56] as reference value, our updated analyses yield residual Gaussian tensions between $1.3\sigma$ and $2.1\sigma$ (down from $\sim 2.5\sigma$ within $\Lambda$CDM).
@@
In addition, for the first time, we have performed a systematic study of the post-phase transition fluid when its equation of state to evolves with time. There are two takeaways from this: (i) The simplifying assumption made in previous work that $w_{\mathrm{NEDE}}=\mathrm{const}$ is still compatible with data. (ii) Nevertheless, data also allows for a significant time evolution that is compatible with a two-component fluid. In particular, such a fluid could be comprised of a stiff ($w=1$) and radiation ($w=1/3$) component with a possible decay channel between both components. The fit indicates that a fraction $r_{f}=0.68^{+0.20}_{-0.13}$ ($68\%$ C.L.) of the latent heat is converted into a stiff fluid with the remaining energy going into a radiation component. While we have merely sketched a few field theoretic scenarios, it should be a priority of future work to come up with detailed and testable descriptions. In particular, each scenario, is expected to affect the perturbation sector in unique ways, which offers the possibility to further decrease the tension.
@@
Assuming future dataset updates do not revert the picture, one way of understanding our results is that the updated datasets indicate that the current form of the Cold NEDE model is not yet complete as a solution to the Hubble tension. This would mean that data are finally able to refine our understanding of early dark energy models, enabling us to explore their unique features required to fully resolve the tension. For Cold NEDE this could involve relaxing assumptions made about the perturbation sector. In particular, it would be interesting to go beyond a tightly coupled barotropic fluid and for example study a more general $k$-dependent sound speed and viscosity parameter (following for example a general fluid approach as in [64]). This phenomenological approach could be informed by one of the different microscopic dissipation scenarios mentioned in this work. Alternatively, it may turn out that there are yet unknown systematic errors affecting either Planck NPIPE data, or the SH0ES measurement, that may go in the direction of improving or relaxing constraints to the model. In any case, Cold NEDE remains so far one of the most promising and natural model for addressing the Hubble tension. Building upon it provides a constructive and promising way forward.

### v2

@@ III.2 Confronting NEDE with NPIPE (and Pantheon+)
We find that Cold NEDE reduces the $Q_{\mathrm{DMAP}}$ tension from $6.3\sigma$ within $\Lambda$CDM to $2.8\sigma$ (Plik/Planck 2018), $3.5\sigma$ (CamSpec/NPIPE), and $3.7\sigma$ (HiLLiPoP/NPIPE). When comparing this with the previous works mentioned above, the residual tension with the newer datasets is larger by up to $1.5\sigma$. However, this comparison has to be taken with a grain of salt: The previous analyses, which relied on Pantheon and Plik/Planck 2018, also gave rise to a weaker $\Lambda$CDM tension. For example, [19] (Plik/Planck 2018+BAO+Pantheon) used a prior on $H_{0}$ that corresponded to a $\Lambda$CDM $Q_{\mathrm{DMAP}}$ tension of $4.7\sigma$, which was reduced to $1.4\sigma$ within Cold NEDE. In other words, the level of relative improvement is similar to what we find here for Plik/Planck 2018 (in both cases, the relaxation of the tension is larger than $3\sigma$). Exchanging Plik/Planck 2018 with CamSpec/NPIPE diminishes this improvement slightly to around $2.5\sigma$. In all cases, we find a significant $\chi^{2}$ improvement of at least 28 units. While compared to previous analyses, the base model remains similarly efficient at reducing the Hubble tension, it cannot fully resolve it anymore for our baseline datasets (provided we accept that the Hubble tension has passed the $Q_{\mathrm{DMAP}}$ threshold of $6\sigma$). The observation that NPIPE data slightly disfavors the NEDE model compared to Plik/Planck 2018 is also supported by the contours in Fig. 1. Comparing the orange (Plik/Planck 2018) with the red (CamSpec/NPIPE) contours, it is apparent that the evidence for NEDE is diminished. To be specific, we find $f_{\mathrm{NEDE}}<8.3\%$ ($95\%$ C.L.) for NPIPE and $f_{\mathrm{NEDE}}<11.9\%$ ($95\%$ C.L.) for Planck 2018.555We stress that this should be understood as a relative comparison as both values are known to be affected by prior volume effects that become sizeable in the $f_{\mathrm{NEDE}}\to 0$ limit [18, 64]. A profile likelihood for $f_{\mathrm{NEDE}}$ can be expected to support a non-vanishing fraction of $f_{\mathrm{NEDE}}$ [61, 24]. Moreover, it should be noted that we are comparing $f_{\mathrm{NEDE}}$ at the time of decay, which is not exactly the same in Plik and NPIPE fits. Specifically, the later decay found for NPIPE might in part explain the difference. This result is reproduced by the blue contours (HiLLiPoP/NPIPE). Overall, we find good agreement between the CamSpec and HiLLiPoP likelihood code. [^6] In addition, we also investigated the internal consistency of the CamSpec likelihood within NEDE by combining our baseline datasets with only the NPIPE TT and the TEEE data. We find consistency between these reduced NPIPE runs and the full one. In particular, we find that polarization data alone has limited constraining power, leaving, for example, $w_{\mathrm{NEDE}}$ and $f_{\mathrm{NEDE}}$ rather unconstrained. The only exception, albeit not significant, concerns the constraint on trigger dark matter $\Omega_{\phi}$, which is $\Omega_{\phi}<0.8\%$ ($95\%$ C.L.) for HiLLiPoP and strengthens to $\Omega_{\phi}<0.5\%$ for CamSpec. In that context, it is important to note that in both cases the model preserves its ability to reduce the value of $S_{8}$, thereby keeping the Gaussian $S_{8}$ tension below $2.1\sigma$ for all analyses performed here (using the weak lensing constraint in [58] as a reference value). Moreover, we find in agreement with the previous analysis in [19] that including the SH0ES prior further reduces the tension to below $2\sigma$. This trend is particularly pronounced in the case of HiLLiPoP.
@@
Determining the origin of the difference between both datasets is nontrivial. To obtain more insight into this, in App. A we construct likelihood curves for both datasets and compare them. Interestingly we observe a bi-modality feature in the NPIPE data, which is absent for Planck 2018.
For the first time, we also tested the Cold NEDE model against the recently released DESI BAO dataset. To that end, we performed a combined analysis with CamSpec/NPIPE and Pantheon+ (green contour in Fig. 1). Notably, this leads with $2.6\sigma$ to the smallest residual tension in this work. A similar tendency of DESI data to relax constraints on $H_{0}$ has been observed recently for the AxiEDE model [31, 32]. This is further confirmed through the likelihood profiles provided in App. A. Notably, when combining with SH0ES, we find a ${}6\sigma$ evidence for a non-vanishing fraction of NEDE, a significance level that had not been obtained in previous analyses.
In Fig. 2, we compare Cold NEDE (in red) and AxiEDE (in blue). We find that both models are similarly affected by NPIPE data, although the constraints on $f_{\mathrm{NEDE}}$ (and hence $H_{0}$) are somewhat weaker, leading to a marginally smaller residual tension in the case of Cold NEDE (in [26] a residual $3.7\sigma$ tension was reported for CamSpec). This raises the expectation that newer CMB datasets will be increasingly capable of distinguishing between the two models (see also Refs. [65, 11] for other comparisons). As mentioned before, while both models share similarities on background level, differences arise predominantly on perturbation level. To be specific, the perturbations in AxiEDE correspond to perturbations of a homogeneous scalar field background and are described by the perturbed Klein-Gordon equation, whereas perturbations in Cold NEDE, are perturbations of a barotropic fluid with an amplitude that, on sub-horizon-scales, is controlled by the momentum-dependent perturbations of a subdominant scalar field evaluated at $\tau_{*}$.
@@ III.3 Constraining a Time-Dependent Equation of State
Figure 3: Left: Comparison of the 2D posterior distributions obtained for the extended NEDE model with a time-dependent equation of state using either the CamSpec/NPIPE (red) or the Plik/Planck 2018 (blue) likelihood (alongside our baseline datasets and the SH0ES prior on $M_{b}$). The green curve shows the prediction for the two-fluid mixture in (17) and (18) (stiff fluid and radiation). The gray lines in the $w$ vs $dw/d\ln a$ plane show the case when the stiff fluid decays into radiation, where the dashed, dash-dotted and dotted lines correspond to $Q_{*}=1/2$, $Q_{*}=1$ and $Q_{*}=2$, respectively. Right: The distribution for the first two derivatives of the equation of state parameter restricted to $w_{\mathrm{NEDE}}(a_{*})=5/6$, demonstrating partial insensitivity to the value of the second derivative (in the case of CamSpec/NPIPE). The black line corresponds to $y=0$. In the region below it, $w$ would turn negative at a later time and, hence, that parameter region is left unexplored.
@@
In the $w$ vs $dw/d\ln a$ plane of Fig. 3, we observe an approximate degeneracy between these two parameters. A larger initial value of $w$ can be compensated by a more negative derivative. A similar degeneracy is observed in the other two panels that depict the $w$ vs $d^{2}w/d(\ln a)^{2}$ and the $dw/d\ln a$ vs $d^{2}w/d(\ln a)^{2}$ planes. There, however, one finds that the width of the degeneracy band increases towards large values of $w$. This feature reflects the fact that in that parameter region, the second derivative is less important for the parametrization of $w(a)$ and does not impact cosmological observables much. At the same time, there is a relatively rigid bound from below corresponding to $d^{2}w/d(\ln a)^{2}\gtrsim 0$ for both data set combinations. This, however, is due to the fact that in our simplified parametrization (10), that parameter region would lead to negative $w(a)$ at late times which is excluded by construction. This is demonstrated in the right panel of Fig. 3 for the fiducial value $w_{\mathrm{NEDE}}(a_{*})=5/6$. In the region below the black line, $w$ turns negative before today. We reiterate that this is an artifact of the truncation in (10). In particular, including higher-order terms is expected to relax that bound. This, however, is not expected to change the results for $dw/d\ln a$, as data can only constrain $w(a)$ over a limited redshift range close to $z_{*}$.
@@
Let us emphasize that the purpose of the extended model is to get an understanding of what fluid features are compatible with data. Ultimately, the details of the fluid, including the time-dependence of its equation of state parameter, will be determined by the details of the underlying model (for example the decay channels of the tunneling field as discussed in Sec. II.3). After all, from a pure data fitting perspective, the mild improvement in $\chi_{\rm min}^{2}$ would not justify adding two more parameters. However, for the purpose of this analysis, this result simply shows that the constraining power current datasets have over the fluid details is still limited, and thus the assumption made in previous works that $w_{\mathrm{NEDE}}$ can be treated as a constant appears justified. We nevertheless evaluated the AIC criterion to provide a more quantitative measure of the overall fit quality in relation to the number of model parameters in Tab 4. While $\Delta{\rm AIC}$ is negative in all cases where the SH0ES prior is included, it is indeed smaller for the base model.
The green curves show the two-fluid constraints (17) and (18) for a mixture with $w_{1}=1$ and $w_{2}=1/3$. Each point on the curve corresponds to a certain choice of the initial ratio $r_{f}$ defined in (16). These results suggests that the mixture model is consistent with both dataset combinations when the two components initially have similar energy densities or the stiff component dominates. For example, for the case of the Plik/ Planck 2018 likelihood, the preferred values of $w_{*}$ translate into the fluid ratio $r_{f}=0.68^{+0.20}_{-0.13}$ ($68\%$ C.L.), according to (16). Note that the case of $r_{f}\ll 1$ (the fluid consisting mostly of radiation at $a_{*}$) is disfavored by data as can be inferred from the $w$ vs $dw/d\ln a$ plane. This is despite the fact that a small second derivative with $w_{*}\approx 1/3$ can be consistent with data as can be seen from the $w$ vs $d^{2}w/d\ln a^{2}$ plane below. However this would require $dw/d\ln a$ to be positive, which is incompatible with the simple mixture model). On the other hand, values of $w_{*}\approx 0.8$ (or $r_{f}\approx 0.75$) are consistent with data according to the $w$ vs $dw/d\ln a$ plane. [^7] It might appear that this in slight tension with the contours in the $w$ vs $d^{2}w/d\ln a^{2}$ plane. This is however only a technical limitation due to the lower bound $w(a)>0$ that we imposed in our analysis. To explore this parameter space one would need to include higher derivative terms.
@@
The gray curves in the figure show the case when the stiff fluid decays into the radiation component with a decay rate controlled by $Q$, according to (20). The dashed, dash-dotted and dotted lines correspond to $Q_{*}=1/2$, $Q_{*}=1$ and $Q_{*}=2$, respectively. For these order unity values, a sizeable fraction of the stiff fluid decays within one Hubble time. In particular, we see that the decay scenarios are compatible with data as long as the decay rate satisfies $Q*\lesssim 2$. In particular, data appears to be compatible with a scenario where $Q_{*}=2$ and $r_{f}=1$, i.e., all of the latent heat is initially converted into a stiff fluid before it decays into radiation rather quickly. The gray curves are not shown in the $d^{2}w/dlna^{2}$ planes because they also depend on the time-dependence of the decay rate, i.e. on the parameter $q_{*}$ according to (22) (making these contours less predictive for the decay scenario).
@@
There are two takeaways from this analysis. First, the time-evolving equation of state is only (mildly) favored for the Planck 2018 data, while it is less constrained for NPIPE data. For both NPIPE and Planck 2018 data, there is a clear (and expected) degeneracy between $dw/d(\ln a)$ and $w_{*}$. Moreover, both datasets are fully compatible with the assumption $dw/d(\ln a)=0$ made in the base model, and accordingly the data fit only gives a very weak evidence for the extension. Second, in both cases, data appears to be compatible with a two-fluid system with $w_{1}=1$ and $w_{2}=1/3$, including the possibility of an energy transfer from the stiff to the radiation fluid. This motivates a more detailed study where both fluid components including their perturbations are evolved independently, without relying on the truncation.
@@ III.4 A First Test of the Thawing Assumption
Remaining agnostic about the underlying model, here, we treat $H_{*}/m$ as a free parameter and fit it to both the CamSpec/NPIPE and Plik/Planck 2018 likelihoods (together with our baseline datasets and the SH0ES prior on $M_{b}$). To be precise, we consider the model with parameters $\log_{10}(z_{*})$, $f_{\mathrm{NEDE}}$, $w_{*}$ ($=\mathrm{const}$), and $H_{*}/m$ (while we set $\Omega_{\phi}\ll 1$ for simplicity). Note that on the background level, for a fixed $z_{*}$, this parameter controls the value of $m$, which enters the equations of motion (4) for the trigger field. Since the trigger field is usually very subdominant, changing its mass has negligible effect on the cosmic evolution. At the perturbation level, it however affects the matching equations (7) through their dependence on $\delta\phi$ and $\bar{\phi}^{\prime}$. In other words, the perturbations in the trigger field at the time of the phase transition control the initial amplitude of perturbations in the decaying NEDE fluid. We see that the matching equation diverges for $\bar{\phi}^{\prime}\to 0$, which during radiation domination imposes the lower bound $H_{*}/m>0.12$. On the other hand, for $H_{*}/m\gg 1$, we enter the slow-roll regime, where the trigger is assumed to be inactive. We impose the flat prior $10.18<H_{*}/m<$ 1. The lower bound corresponds to the first zero of $\bar{\phi}$ after it has dropped out of slow-roll. In a specific trigger scenario such as Cold NEDE, we thus expect the phase transition to occur earlier, i.e., for larger values of $H_{*}/m$. The upper bound corresponds to a situation where the transition is triggered while $\phi$ is still marginally within the slow-roll regime (for even higher values the trigger perturbations are too suppressed for data to be sensitive to differences between different values of $H_{*}/m$).
@@ IV Conclusion
The Cold NEDE model introduces an early dark energy component that decays in a phase transition triggered by an ultralight scalar field. In the past, the model has been found to reduce both the $H_{0}$ and $S_{8}$ tensions below the $2\sigma$ level. It admits a simple field theoretic realization in terms of a first-order vacuum phase transition that occurs in the dark sector before matter-radiation equality. Cold NEDE features some important distinctions from the axiEDE model, which were highlighted in this work. AxiEDE involves only one scalar field which undergoes a smooth crossover transition and has a negligible relic abundance today. Instead, in Cold NEDE there is a first-order vacuum phase transition that is triggered by a scalar field when it thaws around matter-radiation equality. Unlike the scalar field in axiEDE, the trigger field in NEDE does not participate in the energy injection around the recombination time itself. It however seeds the perturbations of the fluid that is used to describe the post-phase transition state. At the same time, the oscillations of the trigger field can have a non-negligible energy today and act as a small fuzzy dark matter component. As a consequence, in contrast to axiEDE and other EDE type models, including the SH0ES prior lowers the value of $S_{8}$ in Cold NEDE [19].
@@
One of the benefits of the Cold NEDE model is its predictivity as it directly affects the physics of waves in the primordial plasma at scales that can be observed in the CMB and matter power spectra. This happens at the background level through a change of the sound horizon and at the perturbation level through the gravitational impact of dark sector perturbations associated with the decaying early dark energy fluid. Nevertheless, the model comes with freedom in how to realize the dissipation of the early dark energy component after the phase transition, which so far has been modeled in terms of an ideal fluid with a constant equation of state parameter $w_{\mathrm{NEDE}}$. At the same time, new CMB, LSS and supernovae observations offer the prospect of further constraining early dark energy models and, in the case of Cold NEDE, this should enable us to further constrain $w_{\mathrm{NEDE}}$ after the phase transition, providing a way to further narrow down viable microscopic scenarios.
@@
In line with this overarching goal, in this work, we have analyzed the cold NEDE model and some of its physically well motivated extensions in light of updated CMB and supernovae data. To be precise, the $-\ell$ Planck 2018 polarization and temperature data and the Pantheon supernovae data were replaced by Planck NPIPE (using the CamSpec and HiLLiPoP likelihoods) and Pantheon+ data, respectively. While we find that the model’s ability to resolve the Hubble tension is somewhat diminished with the inclusion of the new datasets, it still reduces the $Q_{\mathrm{DMAP}}$ tension considerably from $6.3\sigma$ within $\Lambda$CDM to $3.5\sigma$ when using the CamSpec likelihood code for the NPIPE data (and $3.7\sigma$ for HiLLiPoP). This outcome is compatible with recent findings for the AxiEDE model, where a residual $3.7\sigma$ tension was reported [26]. It is important to note that the improvement over $\Lambda$CDM is almost unchanged compared to earlier works when it comes to relieving the Hubble tension, and the increased residual tension is mostly due to $\Lambda$CDM itself showing a stronger tension when using NPIPE and Pantheon+ data. Moreover, we performed a first test of the Cold NEDE model against the recently released DESI BAO data. We find that the tensions is significantly reduced to $2.6\sigma$ when combining DESI with Pantheon+ and NPIPE data (also improving on the recent axiEDE result of $3.1\sigma$ in [32]). A complementary analysis relying on profile likelihoods confirms the above results.
In this work we find that the Cold NEDE model’s ability to help with the $S_{8}$ tension is maintained in the presence of the new datasets. For concreteness, when using the weak lensing constraint from [58] as reference value, our updated analyses yield residual Gaussian tensions between $1.3\sigma$ and $2.1\sigma$ (down from $\sim 2.5\sigma$ within $\Lambda$CDM).
@@
In addition, for the first time, we have performed a systematic study of the post-phase transition fluid when its equation of state evolves with time. There are two takeaways from this: (i) The simplifying assumption made in previous work that $w_{\mathrm{NEDE}}=\mathrm{const}$ is still compatible with data. (ii) Nevertheless, data also allows for a significant time evolution that is compatible with a two-component fluid. In particular, such a fluid could be comprised of a stiff ($w=1$) and radiation ($w=1/3$) component with a possible decay channel between both components. The fit indicates that a fraction $r_{f}=0.68^{+0.20}_{-0.13}$ ($68\%$ C.L.) of the latent heat is converted into a stiff fluid with the remaining energy going into a radiation component. While we have merely sketched a few field theoretic scenarios, it should be a priority of future work to come up with detailed and testable descriptions. In particular, each scenario is expected to affect the perturbation sector in unique ways, which offers the possibility to further decrease the tension. As another straightforward extension, we freed up the parameter combination $H(z_{*})/m$, which determines by how much the trigger field has left the slow-roll regime when it triggers the phase transition. Interestingly, we find first evidence that data prefers a value that is close to the theoretical prediction of $H(z_{*})/m\approx 0.2$ obtained in the tunneling scenario, which we reviewed in Sec. II.3.
@@
Assuming future dataset updates do not revert the picture, one way of understanding our results is that the updated datasets indicate that the current form of the Cold NEDE model is not yet complete as a solution to the Hubble tension. This would mean that data are finally able to refine our understanding of early dark energy models, enabling us to explore their unique features required to fully resolve the tension. For Cold NEDE this could involve relaxing assumptions made about the perturbation sector. In particular, it would be interesting to go beyond a tightly coupled barotropic fluid and for example study a more general $k$-dependent sound speed and viscosity parameter (following for example a general fluid approach as in [66]). This phenomenological approach could be informed by one of the different microscopic dissipation scenarios mentioned in this work. Alternatively, it may turn out that there are yet unknown systematic errors affecting either Planck NPIPE data, or the SH0ES measurement, that may go in the direction of improving or relaxing constraints to the model. In fact, the first results we obtained from including recent DESI BAO data relax the constrain significantly, and future BAO data will be thus be of utmost importance to test the model. In any case, Cold NEDE remains so far one of the most promising and natural model for addressing the Hubble tension. Building upon it provides a constructive and promising way forward.